### v1

# DLCR: A Generative Data Expansion Framework via Diffusion for Clothes-Changing Person Re-ID

Abstract
With the recent exhibited strength of generative diffusion models, an open research question is if images generated by these models can be used to learn better visual representations. While this generative data expansion may suffice for easier visual tasks, we explore its efficacy on a more difficult discriminative task: clothes-changing person re-identification (CC-ReID). CC-ReID aims to match people appearing in non-overlapping cameras, even when they change their clothes across cameras. Not only are current CC-ReID models constrained by the limited diversity of clothing in current CC-ReID datasets, but generating additional data that retains important personal features for accurate identification is a current challenge. To address this issue we propose DLCR, a novel data expansion framework that leverages pre-trained diffusion and large language models (LLMs) to accurately generate diverse images of individuals in varied attire. We generate additional data for five benchmark CC-ReID datasets (PRCC, CCVID, LaST, VC-Clothes, and LTCC) and increase their clothing diversity by $10$x, totaling over $2.1$M images generated. DLCR employs diffusion-based text-guided inpainting, conditioned on clothing prompts constructed using LLMs, to generate synthetic data that only modifies a subject’s clothes while preserving their personally identifiable features. With this massive increase in data, we introduce two novel strategies - progressive learning and test-time prediction refinement - that respectively reduce training time and further boosts CC-ReID performance. We validate our method through extensive ablations and experiments and show massive improvements when training previous CC-ReID methods with our generated data. On the PRCC dataset, we obtain a large top-1 accuracy improvement of $11.3\%$ by training CAL, a previous state of the art (SOTA) method, with DLCR-generated data. We publicly release our code and generated data for each dataset here: https://github.com/CroitoruAlin/dlcr.
1 Introduction
A majority of computer vision research can be categorized as discriminative methods, wherein improved or novel model architectures are proposed and shown to learn better visual representations than their predecessors. These methods operate under the assumption that their training datasets are sufficiently large and diverse, with the limiting factor being the model’s ability to learn the desired objective. However, when there is a lack in diversity in the commonly used datasets for some task, the limiting factor in this case becomes the training data itself. Collecting additional in-domain data can be both difficult and expensive, and only recently have generative models become viable enough to generate additional, high-fidelity training data (which we refer to as generative data expansion). Generative data expansion has been previously explored for supervised or self-supervised learning on easier discriminative tasks, such as ImageNet classification [3, 50]. While there are other unexplored, more difficult discriminative visual tasks that may benefit even more so from generative data expansion, the difficulty lies in generating accurate, in-domain data. In this work, we tackle this problem in the context of CC-ReID, which is difficult in nature and, due to its relative infancy, a majority of datasets suffer from only containing a small number ($2-5$) of clothing outfits per subject [52, 58, 39, 15, 45].
In Person Re-identification [60, 61], the primary objective is to match an individual across images or videos captured by cameras at different times and locations. This task becomes notably challenging due to disparities in the body poses, lighting, backgrounds, viewpoints, and potential occlusions present across cameras. Traditional Re-ID approaches [5, 8, 38], also known as short-term Re-ID, assume that individuals are observed for relatively brief durations and that their attire remains constant throughout videos. However, these assumptions often fall short of practical real-world scenarios. For instance, people’s clothes can change when captured across different camera feeds at different times and days. Consequently, a new branch of Re-ID, known as clothes-changing Re-ID [39, 15, 24, 17], has recently emerged to reliably identify people even when their clothing constantly change over extended periods of time. Recent CC-ReID works commonly make discriminative modifications to traditional Re-ID model architectures and loss formulations to disentangle structural (clothes-independent) and appearance information (clothes-dependent) in the learned feature space. However, these discriminative approaches are inherently constrained by the aforementioned low clothing variations present in current CC-ReID datasets.
Figure 1: Best viewed with zoom. (Left) Comparing alternative data expansion approaches to our method. Simply changing the color of the clothes using our clothing masks preserves the subject’s ID information, but lacks diversity. Using standard diffusion without inpainting can sometimes introduce diversity, but the ID-related information in an image is destroyed. Our method strongly increases clothing diversity, even with specific and difficult prompts, and preserves the subject’s ID-related information for effective CC-ReID. (Right) DLCR significantly increases the size of training data for five benchmark CC-ReID datasets, generating over $2.1$M diverse, clothes-changed images.
To address these issues, in this work we propose an effective data expansion framework for CC-ReID (named DLCR) that increases the number of clothing variations in CC-ReID datasets by intelligently harnessing generative AI and foundation models (Diffusion and large Language models for CC-ReID). DLCR consists of two stages: (1) Data Generation and (2) Re-ID Training and Prediction Refinement. In the first stage, our objective is to enrich a given CC-ReID training dataset by generating additional samples where clothing items are artificially changed through inpainting. With more diverse training images of a particular person in different clothes, a CC-ReID model will learn better clothes-invariant person features during training. To achieve this objective, we utilize diffusion models [20, 42, 6, 36], which are capable of generating realistic images and inpainting for image editing. However, diffusion models often struggle with preserving the intricate details of a person’s identity during image generation (Fig. 1). We overcome this by leveraging a human parsing method [26] to produce binary masks that mark only the clothing regions in an image. We then use this binary mask to retain ID-specific portions in an image during diffusion inpainting, such as the face, hair, and body shape, thereby only augmenting the subject’s clothes. Furthermore, existing CC-ReID datasets only use scalar values for each unique clothes ID and lack textual descriptions of specific clothing items, preventing the use of text-guided diffusion models for data generation. Thus, to overcome this issue, we generate explicit clothing descriptions by jointly leveraging a large visual-language model (LLaVA [32]) and a large language model (LLaMA [51]).
In the second stage of DLCR, we propose two effective strategies in which our generated data can be sophisticatedly utilized during CC-ReID training and testing. Firstly, we train various CC-ReID models using our generated data from stage 1 and show how our enriched training data improves performance in clothes-changing settings. To accommodate both the large increase in data and the higher clothing variety of the generated samples, we employ a progressive learning strategy during training, where we gradually introduce the generated clothing variations of a person. This allows for training on $10$x more data without drastically increasing training time, as well as further assists the model in learning better clothes-invariant person features. During inference, we apply our prediction refinement strategy which is specifically tailored for a CC-ReID retrieval task. We create different variations of a given query by artificially changing the subject’s attire, then ensemble their retrieved similarity scores from the gallery to obtain a refined final prediction.
We summarize our contributions as follows:
  • DLCR is the first to implement a text-guided diffusion approach, in conjunction with foundational language models, to synthesize multiple images of a person with different clothes in a CC-ReID dataset (Sec. 3.1).
  • We propose a progressive learning strategy during training that judiciously combines our generated data with the original training set, improving performance and heavily reducing training time (Sec. 3.2.1).
  • We employ a novel prediction refinement strategy at test-time, where variants of a given query image are created by changing the subject’s clothes, and their retrieval scores are ensembled to further boost performance (Sec. 3.2.2).
  • We demonstrate the effectiveness of our method on four benchmark CC-ReID datasets (Sec. 4.1). Our method also yields a significant performance boost when existing approaches (both standard Re-ID and CC-ReID) utilize our generated data during training (Sec. 4.2).
  • We publicly release our code and generated CC-ReID data, totaling over $2.1$M generated images, for future use.
2 Related Works
In standard Re-ID, clothing is considered a static attribute, which leads standard Re-ID models to erroneously couple clothing items with specific subjects. CC-ReID is a more difficult paradigm that requires models to intrinsically learn person-specific features that are robust to covariate information, especially clothing items. To achieve this robustness, both discriminative [19, 16, 21, 58, 28, 11, 62, 31, 56, 18] and generative [35, 55, 13, 34, 63, 57] approaches for CC-ReID have been explored. Discriminative approaches focus on deliberate model architectures [30, 15] and loss formulations [19, 16, 21], while generative methods aim to synthesize data that increases the quality or diversity of an existing training set. The common goal of these methods is to enforce the learning of clothes-invariant person features, thereby improving CC-ReID performance on datasets in which subjects are captured in multiple different clothing outfits.
Discriminative methods: A standard approach to learning clothes-invariant features is to use some form of identification loss paired with a metric learning loss, i.e. contrastive or triplet loss [19, 16, 21, 58]. Alternatively, some methods specifically aim to disentangle the appearance information from the person-relevant information in RGB images [28, 11, 62] for CC-ReID. Tangential to metric learning, other works have also explored the use of adversarial learning for CC-ReID [30, 15]. Discriminative methods are not limited to loss formulations, as novel architectural designs have also been proposed to accommodate additional clothes-agnostic cues, such as face extraction [2] and 3D body shape reconstruction [31]. Similarly, auxiliary input modalities to RGB, such as human pose [54, 48], depth [56, 18], and gait [24, 27] have been proposed as multi-modal discriminative approaches to CC-ReID. Separate from input modalities, other works directly augment the learned feature space for better semantic control [17] or robustness to covariate information [46]. However, these discriminative methods are constrained by the inherent limitations of the datasets they are trained on. For example, an insufficient number of clothing outfits in a dataset can be detrimental to metric learning-based CC-ReID models [17]. Thus, our work proposes a generative method that directly improves the clothing diversity in a dataset, better equipping any downstream CC-ReID model to learn clothes-invariant features.
Generative methods: With the aforementioned lack of such diverse datasets, the previously dominant approach has been to use Generative Adversarial Networks (GANs) [14] to synthesize additional training data for specific learning objectives, such as pose-invariance [35, 55, 13], clothes-invariance [34, 63, 57], and improving general performance [64, 29, 23]. However, these GAN-based works also suffer from constraints such as limited diversity/mode collapse [63, 33] and the limited capability of only transferring existing styles from one subject to another [9, 65, 66]. Most important to note, GANs are dataset-specific, meaning they require separate training on each dataset before data generation. Furthermore, one of the most successful GAN-based works [63] does not directly use its generated data during training, but rather transfers its dataset-specific generator encoder to the discriminator to produce better clothes-invariant person features, reducing generalizability. On the other hand, diffusion models [36] have recently overtaken GANs in a multitude of other vision domains [10, 6, 43, 44] as a superior class of generative models, yet have not been applied to the CC-ReID domain. Furthermore, the previously discussed limitations of GANs can be overcome by large-scale, pre-trained diffusion models which have shown impressive zero-shot generation abilities on a multitude of out-of-domain tasks [40, 42, 37]. DLCR is the first generative approach to harness the benefits of diffusion models for CC-ReID, and consequently, uses the same generator for every dataset. In a novel fashion, we leverage a text-conditioned diffusion inpainting model to synthesize completely new clothing outfits onto any subject in a dataset, separating the process between data generation and discriminative CC-ReID training.
3 Proposed Approach
DLCR is a generic data expansion method that can be applied to any CC-ReID dataset, wherein the given training data is enriched to improve data diversity for better learning of clothes-invariant features. Our method contains two stages: Data Generation (Sec. 2) and Re-ID Training and Prediction Refinement (Sec. 3.2). In stage 1, DLCR utilizes semantic human masks, a visual language model (LLaVA), a large language model (LLaMA), and text-conditioned diffusion inpainting to controllably change a subject’s clothes in an image. In stage 2, we leverage our generated data, in conjunction with our progressive learning strategy, to train CC-ReID models with $10$x more data with negligible impacts to training time. We also propose a test-time prediction refinement strategy specifically tailored for CC-ReID retrieval. Before explaining the method in detail, we provide a brief background on diffusion models.
Diffusion models: are a class of likelihood-based generative models that learn a data distribution by parameterizing a Markov chain through the use of forward and backward processes. Given a data sample, $x_{0}$ $\sim q(x_{0})$, the forward process of diffusion iteratively adds Gaussian noise, $\epsilon$, to $x_{0}$ for $T$ timesteps until $x_{T}$ is almost a sample from a standard Gaussian distribution. Then, the backward process leverages a neural network to learn the reverse distribution $q(x_{t-1}|x_{t})$. Thus, a novel data point from the original data distribution, $q(x_{0})$, can be simply generated by sampling $x_{T}$ from $\mathcal{N}(\textbf{0},\textbf{I})$ and applying the Markov reverse steps until $x_{0}$ is reached.
Forward Process: At each timestep, Gaussian noise is added with a timestep dependent variance $\beta_{t}$ to obtain $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\textbf{I})$. We can also efficiently sample $x_{t}$ in a single step [20] as follows:
$x_{t}\sim q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha_{t}})\textbf{I}),$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$.
Reverse Process: The reverse distribution $q(x_{t-1}|x_{t})$ is approximated with a parameterized model $p_{\theta}$. With small enough chosen values of $\beta_{t}$, $q(x_{t-1}|x_{t})$ will be Gaussian, so a neural network with parameters $\theta$ can be trained to learn the mean $\boldsymbol{\mu}_{\theta}(x_{t},t)$ and variance $\mathbf{\Sigma}_{\theta}(x_{t},t)$ at each timestep $t$:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\boldsymbol{\mu}_{\theta}(x_{t},t),\mathbf{\Sigma}_{\theta}(x_{t},t)).$ (2)
Ho et al. [20] simplifies this objective by expressing $\boldsymbol{\mu}_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\left\lVert x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right\rVert\right),$ and propose to learn to predict $\epsilon_{\theta}(x_{t},t)$ using the loss function $\mathcal{L}_{simple}=E_{t,x_{t},\epsilon}(\lVert\epsilon-\epsilon_{\theta}(x_{t},t)\rVert^{2})$, where $\mathbf{\Sigma}_{\theta}(x_{t},t)$ is fixed to $\beta_{t}\mathbf{I}$. Our method uses text-conditioned diffusion, where a text prompt is used to guide the generation at every timestep, starting from timestep $t=T$ to $t=0$. The reverse process can simply be modified to incorporate a condition $y$ as follows:
$p_{\theta}(x_{t-1}|x_{t},y)=\mathcal{N}(x_{t-1};\boldsymbol{\mu}_{\theta}(x_{t},t,y),\beta_{t}\mathbf{I}).\vspace{-0.15in}$ (3)
3.1 Stage 1: Data Generation
In CC-ReID, the training dataset $D_{train}=\{(x_{i},s_{i},c_{i})\}_{i=1}^{M}$, is a collection of $M$ samples where an $i^{th}$ sample is a triplet consisting of an RGB image $x_{i}$, subject (person) ID $s_{i}$, and clothes ID $c_{i}$. Generating images of a particular subject, $s_{i}$, in different clothing outfits can enrich the training data by improving its diversity. Specifically, in this stage, we generate multiple variants of a training image, $x_{i}$, by artificially inpainting the clothes of subject $s_{i}$. We achieve this objective by using a text-conditioned diffusion inpainting model (Stable Diffusion Inpainting [42]), where text prompts describe the outfits to be generated in the clothing regions specified by a human-parsed semantic mask. These components (also shown in Fig. 2) are described below.
Figure 2: Overview of DLCR data generation (stage 1). Given a training image $x_{i}$, we apply an ID-preserving mask, $m_{i}$, where only the clothing regions are marked for inpainting. A target clothing ID, $c$, is randomly selected from the training dataset and a textual description of the clothing items belonging to $c$ is extracted. Specifically, we use LLaVA to extract frame-wise descriptions for every image in $X^{c}$, and use LLaMA to obtain a summarized description $R^{c}$. This description is used as a prompt for diffusion inpainting to artificially change a subject’s clothing. Our ID-preserving diffusion inpainting model synthesizes quality generated data $\hat{x}_{0}$ for improved downstream CC-ReID training.
Semantic Mask & Text Extraction: While generating different images of a subject, it is important to preserve the person-relevant information. However, diffusion models struggle to retain this information due to the varied and complex nature of human faces and body shapes. Thus, we apply a semantic human parsing method [26] to each image, $x_{i}$, to obtain an ID-preserving binary mask, $m_{i}$, where only the upper-clothes, lower-clothes, and footwear regions are marked for inpainting. This process is computationally negligible, with throughput as high as 60 masks per second. This allows the person ($s_{i}$) features, such as the face, hair, and general body structure, to be preserved in the generated samples, even after clothes inpainting; an especially critical aspect for enriching CC-ReID data. In general, the set of images belonging to a clothes ID, $c$, can be represented by $X^{c}=\{x^{c}_{i}\}_{i=1}^{n}$. After applying ID-preserving semantic masks to these images, we obtain a set of masked images $\bar{X}^{c}=\{x^{c}_{i}\circ m_{i}\}_{i=1}^{n}$. While CC-ReID datasets provide clothes IDs, they are usually simple scalar values that do not contain descriptive information as to what specific clothing items are present in an image. To extract these additional clothing descriptions for each clothes ID in $D_{train}$, we utilize LLaVA [32] and LLaMA [51] (see Fig. 2). A naive approach would be to simply generate descriptions of clothing items using an LLM, e.g. LLaMA, or to create random clothing descriptions. While this is likely to increase the diversity of the dataset, and consequently, the generalization capacity of the downstream CC-ReID models, it does not alleviate dataset-specific biases. We use LLaVA in order to obtain descriptions of clothing items that are present in the dataset, aiming to reduce the dataset-specific bias (see supplementary). This forces the CC-ReID models to focus on identity features, and ignore clothing features.
First, we prompt LLaVA to obtain the information of the clothing items for the upper and lower body, as well as the footwear a person is wearing in each image $x_{i}^{c}\in X^{c}$. We denote the set of text responses obtained from LLaVA as $r^{c}=\{r^{c}_{i}\}_{i=1}^{n}$, where $r^{c}_{i}$ is the clothes description of image $x_{i}^{c}$. However, the subject’s visibility can vary across images of a particular clothes ID, $c$, due to changes in occlusions, lighting, or viewpoints. This can lead LLaVA to occasionally respond with missing or incorrect clothing items on certain images of $X^{c}$. To mitigate this issue, we pass the image-based responses, $r^{c}$, as input to LLaMA and prompt it to summarize them into a single clothing description. Consequently, we obtain a holistic clothes description, $R^{c}$, for a particular clothes ID $c$. Through this summarization, LLaMA helps to produce accurate clothing descriptions and overcomes the issue of missing clothing items. We further post-process $R^{c}$ if there are multiple clothing descriptions obtained from LLaMA for a particular body part. We keep the most frequently occurring descriptions across the images of $c$. We do not train/fine-tune the LLMs as we only use them in an inference setting with an average throughput of roughly 4 samples per second. For more details regarding our use of LLMs and their impact on the final results, please refer to the supplementary.
Diffusion-Based Clothes Inpainting: A text-conditioned diffusion inpainting model, paired with our extracted clothing descriptions and semantic masks, can be used to controllably generate diverse, synthetic samples. To this end, we propose, for the first time, the use of a diffusion inpainting model for CC-ReID. We first construct a set $C=\{c_{i}\}_{i=1}^{L}$ containing all the unique clothes IDs in the training data $D_{train}$. For each training image $x_{i}$ with subject ID $s_{i}$ and clothes ID $c_{i}$, we uniformly sample $K$ different clothes IDs, $c_{j}\sim\mathcal{U}(C),\forall j\in\{1,...,K\}$ where $j\neq i$. For a sampled clothes ID $c_{j}$, we obtain its holistic clothes description $R^{c_{j}}$ by leveraging LLaVA and LLaMA (Sec. 2). For a given image $x_{i}$, we get the ID-preserved masked image $\bar{x}_{i}$ by applying its corresponding human-parsed semantic mask, $m_{i}$. To generate $K$ diverse images of the same subject (person $s_{i}$) corresponding to the sampled $K$ clothes IDs, we utilize the extracted clothes descriptions $\{R^{c_{j}}\}_{j=1}^{K}$ and masked image $\bar{x}_{i}$ as inputs to a pretrained Stable Diffusion model [42] for ID-preserving inpainting. The masked image $\bar{x}_{i}$ is used in the forward diffusion process (Eq. (1)), to preserve ID-related information. The reverse denoising process (Eq. (3)) is conditioned on the prompt $R^{c_{j}}$ to generate the clothes-changed images $\{\hat{x}_{ij}\}_{j=1}^{K}$ of subject $s_{i}$ in $x_{i}$, as follows:
$\hat{x}_{ij}\sim p_{\theta}({x}_{0}|\bar{x}_{i},R^{c_{j}}),\forall j\in\{1,...,K\}.$ (4)
We do not train/fine-tune the pretrained diffusion model, but solely use it in an inference setting, thus keeping computational requirements low with a throughput of roughly one generated image per second. Thus, we construct a set of generated data $G_{train}=\{(\hat{x}_{ij},s_{i},c_{ij})_{j=1}^{K}\}_{i=1}^{M}$ for each aforementioned CC-ReID dataset which can be directly utilized to enrich the original training data set, $D_{train}$. It is this data, containing over $2.1$M generated images in total across multiple CC-ReID datasets, that we release for public use as summarized in Table 1.
Table 1: Dataset statistics for each CC-ReID dataset with and without DLCR. (*) We do not publicly release our generated data for LTCC due to its licensing restrictions.
  Dataset Training Images
Original Generated
  PRCC $17,896$ $178,960$
LTCC* $9,576$* $95,760$*
CCVID $118,613$ $1,186,130$
VC-Clothes $9,449$ $94,490$
LaST $71,248$ $712,480$
  Total - $2,267,820$
3.2 Stage 2: Re-ID Training and Refinement
Our generated training set $G_{train}$ can be used to enrich the original training data $D_{train}$, improving the discriminative capabilities of any Re-ID model. One simple way to utilize $G_{train}$ is to concatenate it with ${D_{train}}$ into one dataset, then use this merged dataset to train a Re-ID model. However, the model struggles to effectively utilize $G_{train}$ when all the clothes-changed variations are introduced at once during training (discussed in supplementary). Moreover, the training time of the CC-ReID models would be heavily impacted with the generated training data being $10$x larger in size. In the supplementary, we provide a t-SNE visualization of real vs. generated images showing that the generated data exhibits a higher variance. To gently and more efficiently accommodate CC-ReID models to the diverse distribution of the generated data, we propose a progressive learning approach that sophisticatedly increases the clothing diversity of a subject as training progresses. Furthermore, the novel ability to controllably inpaint realistic clothes using our method can be extended even further than just during training. Specifically, we propose a prediction refinement strategy that uses our diffusion-based inpainting method as a unique test-time augmentation to provide a more holistic and ensemble-based decision during evaluation.
3.2.1 Progressive Learning Strategy
To mitigate impacts to training time, we gradually introduce $G_{train}$ by injecting our generated data into each training batch in a pairwise fashion. During batch construction, we pair every original training image in the batch with one of its corresponding generated samples, effectively doubling the original batch size. For brevity, consider one image from the original training set $x\in D_{train}$ and its corresponding generated samples $P^{x}=\{\hat{x}_{i}\}_{i=1}^{K}$. A generated image from $P^{x}$ can be chosen as a pair for $x$ through random sampling: $\hat{x}\sim\mathcal{U}(P^{x})$. Due to the large increase in clothing diversity and number of images in $G_{train}$, we show that incrementally introducing the clothes-changed generated images from $P^{x}$ as training progresses yields better results. Moreover, since we generate roughly $10$x more training data for each CC-ReID dataset, we found that our progressive learning strategy enables models to effectively utilize our generated data without major impacts to training time (see supplementary). Specifically, we perform progressive learning by first dividing $P^{x}$ into four distinct, equal partitions $P^{x}=P_{1}+P_{2}+P_{3}+P_{4}$. Then, during the early stages of training, only the generated images $\hat{x}\in P_{1}$ are used for sampling when finding a pair for $x$, where $\hat{x}\sim\mathcal{U}(P_{1})$. As training continues, we periodically increase the sampling range for the generated images by incorporating more partitions. For instance, $x$ is paired with a randomly sampled generated image $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}\})$ after $N$ epochs of training have been completed. After $2N$ completed training epochs, the generated pair image is sampled by $\hat{x}\sim\mathcal{U}(\{P_{1}+P_{2}+P_{3}\})$ until eventually, every generated image in $P^{x}$ is used for sampling during batch construction after $3N$ training epochs: $\hat{x}\sim\mathcal{U}(P^{x})$. In this manner, we increase the difficulty of learning clothes-invariant features by controlling the number of clothes-changed generated images seen during the early stages of training. Therefore, the model gradually learns to ignore more clothing items as training progresses due to the increasing diversity of subject attire.
3.2.2 Prediction Refinement Strategy
While we have only discussed using DLCR for generating CC-ReID training data, we exhibit in this section that we can also leverage DLCR at test-time by employing our inpainting method (Sec. 2) to augment a given query image by changing its clothes. For a given query, we create $l$ augmented query variants. For each query variant, we then retrieve the top-$m$ similarity scores from the gallery and ensemble them to mitigate clothing bias when making a prediction. Consequently, our ensembling method yields a refined similarity score estimate of a subject ID, resulting in more accurate top-1 retrievals. Algorithm 1 describes the detailed steps of our ensembling process. Our algorithm takes a query image, the $l$ augmented versions of the query, the gallery, and the CC-ReID model as input. The first step computes the features of the query ($f^{q}$), augmented query variants $\{\hat{f}^{q}_{i}\}_{i=1}^{l}$ and gallery images $\{\hat{f}^{g}_{i}\}_{i=1}^{n}$ (Line $1$). In step two, for the original query, we retrieve the top-$m$ most similar examples, $I^{m}$, from the gallery (Lines $2$-$3$). In the fourth step, each unique subject $\{s_{i}\in S\}_{i=1}^{m}$ in $I^{m}$ is assigned an initial score. We repeat this retrieval process for each augmented query variant $\{\hat{f}^{q}_{k}\}_{k=1}^{l}$, resulting in their own top-$m$ retrievals from the gallery, $\{I_{k}^{m}\}_{k=1}^{l}$ (Line $5$). For each $I_{k}^{m}$, we update the similarity scores only for the subjects present in $S$ (Line $6$). At the end, we obtain the set of subject IDs $S$, with their corresponding refined similarity scores. Hence, for a given query image, we leverage our diffusion-based inpainting and ensembling algorithm to yield a better clothes-invariant prediction. Algorithm 1 is also extremely computationally efficient, with the entire process having a throughput of $\sim 640$ images/sec with a memory constraint of $<2$ GB.
Input: $x^{q}$: query image, $\{(x^{g}_{i},s_{i})\}_{i=1}^{n}$: labeled gallery images,
$\{\hat{x}^{q}_{i}\}_{i=1}^{l}$: $l$ augmented variants of $x^{q}$, $m$: size of query-gallery retrieval,
$r_{\psi}$: CC-ReID model, $\cos(\cdot,\cdot)$: cosine similarity.
Output: Set $S$ of subject IDs.
/* 1: Extract features */
1 $f^{q}\leftarrow r_{\psi}(x^{q});f^{g}_{i}\leftarrow r_{\psi}(x^{g}_{i}),\forall i\in\{1,...,n\};$ $\hat{f}^{q}_{i}\leftarrow r_{\psi}(\hat{x}^{q}_{i}),\forall i\in\{1,...,l\}$
/* 2: Get top-m similar gallery samples */
2 $I\leftarrow\text{argsort}_{i\in\{1,...,n\}}\left\{\cos(f^{q},f^{g}_{i})\right\}$ //  desc order
3 $I^{m}\leftarrow I\left[:m\right]$
/* 3: Assign similarity scores to subject id in top-m */
4 $S\leftarrow\{s_{i}|i\in I^{m}\}$; $score\left[s\right]\leftarrow\sum_{j\in I^{m}}{\mathbbm{1}_{\{s\}}(s_{j})\frac{\cos(f^{q},f^{g}_{j})}{\max\limits_{k\in I^{m}}\cos(f^{q},f^{g}_{k})}},\forall s\in S$
for $k=1,\dots,l$ do
[…]
       /* 4: Update the similarity scores using the inpainted examples */
       for $s\in S_{k}$ do
             if $s\in S$ then
                  6 $score\left[s\right]\leftarrow score\left[s\right]+\sum_{j\in I_{k}^{m}}{\mathbbm{1}_{\{s\}}(s_{j})\frac{\cos(\hat{f}^{q}_{k},f^{g}_{j})}{\max\limits_{p\in I_{k}^{m}}\cos(\hat{f}^{q}_{k},f^{g}_{p})}}$
             end if
       end for
end for
7return $S$
Algorithm 1 Refinement of Re-ID Predictions
4 Experiments
Implementation details: We use CAL [15] as the baseline CC-ReID model in our experiments as it generally performed the best with our generated data across all datasets. We follow all of their training hyper-parameters unless specified otherwise, but note that our method is compatible with any Re-ID model as discussed in Sec. 3. We use Stable Diffusion [42] as our diffusion inpainting model. We set the number of generated clothing outfits to $K=10$. In progressive learning, we increase the sampling range after every $N=5$ epochs. For our prediction refinement strategy, we set the number of augmented queries $l=6$ and keep the top $m=5$ retrievals. We resize input images to $384\times 192$, and train downstream Re-ID models with an original batch size of $32$ which is doubled to $64$ when paired with generated data, however note that this has a negligible effect on Re-ID training time. All experiments are conducted on a single NVIDIA A100 80GB GPU, but only the data generation stage requires at least one GPU with $\leq$ 48GB of memory, whereas training and deploying the CC-ReID models can be done on smaller consumer GPUs ($\leq$12GB). Additional throughput and computational complexity details are provided in the supplementary.
4.1 Comparison with State-of-the-Art Methods
We report results on the clothes-changing settings of four recent CC-ReID datasets: PRCC [58], LTCC [39], VC-Clothes [52], and CCVID [15]. Following the standard protocol described in each dataset, we use the top-1 accuracy and mean average precision (mAP) as evaluation metrics, and use the official dataset query and gallery splits. We apply stage 1 (Sec. 2) and stage 2 (Sec. 3.2) of DLCR to CAL [15] to obtain our results shown in Table 2. With the help of DLCR-generated data, we achieve new SOTA on 3 out of 4 datasets with CAL, with the largest SOTA improvement of $5.3\%$ in top-1 accuracy on PRCC. Note that while CCFA [17] is the only published model to outperform DLCR on LTCC, their source code is not available, hence we were not able to train CCFA using our generated data. With available code, we hypothesize that DLCR-generated data would most likely improve CCFA and achieve SOTA on LTCC.
Table 2: Comparisons with previous CC-ReID models on the PRCC, LTCC, CCVID, and VC-Clothes datasets in the clothes-changed (CC) setting. DLCR outperforms basically every model across four datasets, with the green text denoting improvements over the CAL baseline. $\dagger$ denotes reproduced results using provided open-source code.
  Dataset Model Venue Top-1 mAP
  PRCC GI-ReID [24] CVPR ’22 37.5 -
AFD-Net [57] IJCAI ’21 42.8 -
RCSANet [22] ICCV ’21 50.2 48.6
3DSL [4] CVPR ’21 51.3 -
FSAM [21] CVPR ’21 54.5 -
CAL (baseline) $\dagger$ [15] CVPR ’22 55.2 55.8
AIM [59] CVPR ’23 57.9 58.3
CCFA1 [17] CVPR ’23 61.2 58.4
Ours - 66.5 (+11.3) 63.0 (+7.2)
  CCVID CAL (baseline) [15] CVPR ’22 81.7 79.6
3DInvarReID [31] ICCV ’23 84.3 81.3
Ours - 88.0 (+6.3) 84.5 (+4.9)
  VC-Clothes GI-ReID [24] CVPR ’22 64.5 57.8
3DSL [4] CVPR ’21 79.9 81.2
FSAM [21] CVPR ’21 78.6 78.9
CAL (baseline) [15] CVPR ’22 85.8 79.8
Ours - 87.1 (+1.3) 81.1 (+1.3)
  LTCC GI-ReID [24] CVPR ’22 26.7 12.7
3DSL [4] CVPR ’21 31.2 14.8
FSAM [21] CVPR ’21 38.5 16.2
CAL (baseline) $\dagger$ [15] CVPR ’22 39.5 18.0
AIM [59] CVPR ’23 40.6 19.1
CCFA [^1] Source code not publicly available to train on our generated data [17] CVPR ’23 45.3 22.1
Ours - 41.3 (+1.8) 19.6 (+1.6)
4.2 Improving Existing Methods using DLCR
Our results in Table 2 are achieved by applying DLCR to CAL since it generally performed the best with our data across all datasets. While both stage 1 (Sec. 2) and stage 2 (Sec. 3.2) of DLCR can be used with any Re-ID model, we show in Table 3 that simply adding our stage 1 generated data during training with any existing method provides significant CC-ReID performance gains. We train three standard Re-ID models (PCB [49], MGN [53], and HPM [12]) and three CC-ReID models (CAL, AIM [59], and GEFF [2]) with our added generated data on the clothes-changing protocol of PRCC. Even when applied to standard Re-ID models that are not specifically tailored to be clothes-agnostic, our generated data boosts CC-ReID top-1 accuracy by nearly $12-30\%$ and mAP by $12-22\%$. Moreover, training CC-ReID models, which are already tailored for clothes-invariance, with our additional generated data provides further top-1 and mAP improvements of $2-8\%$. In the supplementary, we show ablations on each proposed component of DLCR and further improvements when both stages of DLCR are applied.
Table 3: Results on the clothes-changing protocol of PRCC when training various standard Re-ID and CC-ReID models using only stage 1 of DLCR. Our generated data improves performance for any model. $\dagger$ denotes reproduced results using open-source code.
  Re-ID Type Model Top-1 mAP
  Standard PCB [49] 41.8 38.7
PCB + DLCR 53.3 (+11.5) 50.7 (+12.0)
MGN [53] 33.8 35.9
MGN + DLCR 62.5 (+28.7) 57.6 (+21.7)
HPM [12] 40.4 37.2
HPM + DLCR 56.0 (+15.6) 50.9 (+13.7)
  CAL $\dagger$ [15] 55.2 55.8
CAL + DLCR 62.9 (+7.7) 60.9 (+5.1)
Clothes AIM $\dagger$ [59] 55.7 56.3
Changing AIM + DLCR 60.2 (+4.5) 59.0 (+2.7)
GEFF [2] 83.6 64.0
GEFF + DLCR 84.6 (+1.0) 66.0 (+2.0)
4.3 Performance on Out-of-Distribution Data
We show in this section that CAL+DLCR improves performance on out-of-distribution testing data as well. As previously mentioned, CAL introduces architectural changes and loss formulations, such as a clothes-classifier and clothes adversarial loss, to enforce clothes-agnostic learning. However, discriminative changes like these may only enable a model to perform better on data that it was trained on, while poorly generalizing when evaluated on unseen data. To this point, we show results in Table 4 where we take a pre-trained CAL model and a pre-trained CAL+DLCR model on the LTCC dataset, and evaluate them in a zero-shot fashion on other CC-ReID datasets (PRCC, LaST [45], and VC-Clothes). Top-1 accuracy gains ranging from $4\%$ to $8\%$ are seen across all three CC-ReID datasets, with similar gains on mAP, ranging from $\approx 1\%$ to $6\%$. Therefore, DLCR both enhances a model’s performance on a single dataset and even on out-of-distribution testing data.
Table 4: Comparisons between CAL+DLCR and CAL models pre-trained on LTCC when evaluated on other datasets. CAL+DLCR consistently beats CAL when evaluated on other datasets in a zero-shot fashion, exhibiting DLCR’s discriminative benefits extend beyond just the training dataset.
  Model Dataset Top-1 mAP
Training Evaluation
  CAL [15] LTCC PRCC 37.5 35.4
LaST 20.7 3.5
VC-Clothes 5.3 5.6
CAL+DLCR PRCC 45.5 38.8
LaST 25.0 4.2
VC-Clothes 11.8 11.9
5 Conclusion
In this work, we proposed the first use of diffusion for generative data expansion in CC-ReID. Through diffusion inpainting, DLCR is capable of controllably generating completely new clothing items onto a subject while retaining person-specific information. Guided by clothing prompts extracted using foundational language models, DLCR synthesizes additional training data to increase the clothing diversity in a CC-ReID dataset. During CC-ReID training, a progressive learning strategy is applied to efficiently utilize this generated data. To better refine a model’s prediction during inference, we employ a test-time augmentation. We demonstrate that DLCR significantly improves CC-ReID performance and achieves SOTA results across multiple benchmark datasets. Finally, we release all of our generated data for the benchmark datasets used in this paper for future research.